Compute a 95% confidence interval for the population mean, based on the numbers 1, 2, 3, 4, 5, 6, 20. Change the number 20 to 7 and recalculate the confidence interval. Using these results, describe the effect of an outlier (i.e., extreme value) on confidence interval.




1
Expert's answer
2021-12-08T11:05:59-0500

1.


"mean=\bar{x}=\dfrac{1}{n}\sum_ix_i=\dfrac{1}{7}(1+2+3+4+5"

"+6+20)=\dfrac{41}{7}\approx5.857143"

"s^2=\dfrac{1}{n-1}\sum_i(x_i-\bar{x})^2=\dfrac{1}{7-1}((1-\dfrac{41}{7})^2+"

"+(2-\dfrac{41}{7})^2+(3-\dfrac{41}{7})^2+(4-\dfrac{41}{7})^2"

"+(5-\dfrac{41}{7})^2+(6-\dfrac{41}{7})^2+(20-\dfrac{41}{7})^2"

"=\dfrac{12292}{294}=\dfrac{1756}{42}"

"s=\sqrt{s^2}=\sqrt{\dfrac{1756}{42}}\approx6.466028"

The critical value for "\alpha = 0.05" and "df = n-1 = 6" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.446899"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(5.857143-2.446899\times\dfrac{6.466028}{\sqrt{7}},"

"5.857143+2.446899\times\dfrac{6.466028}{\sqrt{7}})"

"=(-0.1229, 11.8372)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "-0.1229 < \mu < 11.8372," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(-0.1229, 11.8372)."


2.


"mean=\bar{x}=\dfrac{1}{n}\sum_ix_i=\dfrac{1}{7}(1+2+3+4+5"

"+6+7)=\dfrac{28}{7}=4"

"s^2=\dfrac{1}{n-1}\sum_i(x_i-\bar{x})^2=\dfrac{1}{7-1}((1-4)^2+"

"+(2-4)^2+(3-4)^2+(4-4)^2"

"+(5-4)^2+(6-4)^2+(7-4)^2"

"=\dfrac{28}{6}=\dfrac{14}{3}"

"s=\sqrt{s^2}=\sqrt{\dfrac{14}{3}}\approx2.160247"

The critical value for "\alpha = 0.05" and "df = n-1 = 6" degrees of freedom is "t_c = z_{1-\alpha/2; n-1} = 2.446899"

The corresponding confidence interval is computed as shown below:


"CI=(\bar{x}-t_c\times\dfrac{s}{\sqrt{n}}, \bar{x}+t_c\times\dfrac{s}{\sqrt{n}})"

"=(4-2.446899\times\dfrac{2.160247}{\sqrt{7}},"

"4+2.446899\times\dfrac{2.160247}{\sqrt{7}})"

"=(2.0021, 5.9979)"

Therefore, based on the data provided, the 95% confidence interval for the population mean is "2.0021 < \mu < 5.9979," which indicates that we are 95% confident that the true population mean "\mu" is contained by the interval "(2.0021, 5.9979)."


3.

The confidence interval is wider in the first case because of the outlier 20.